(CO 1) What is the standard deviation and its meaning given this population of customer ages? 45, 76, 30, 22, 51, 40, 63, 66, 41

16.5 which means that many of the ages will be around this value
17.5 which means that many of the ages will be around the square of this value
16.5 which means that many of the ages will be within this value of the average age
17.5 which means that many of the ages will be within this value of the average age
Answer #1

Sample mean \bar{x} = \sum X / n

= 434 / 9

= 48.2222

Standard deviation = sqrt [ \sum X2 - n \bar{x}2 / n-1 ]

= sqrt [ 23392 - 9 *43.782 / 8 ]

S = 17.5

17.5 Which means that meany of the ages will be within this value of the average age.